A 3 phase, 440 V, 6 pole, 970 rpm, 50 Hz, Y-connected induction motor has the following parameters referred to the stator Rs = 0.2 Ω, Rr一0.15 Ω, xs-x-= 0.4 Ω The stator to rotor turns ratio is 3.5 The motor speed is controled by the Static Scherbius Drive. The drive is designed for a speed range of 30% below the synchronous speed. The maximum value of firing angle is 170. Calculate (i) Turns ratio of the transformer (ii) Torque for a speed of 750 rpm and α = 140 ng angle for half the rated motor torque and a speed of 850 rpm.
Answer #1

LL·490 rat = eo.is 3-5 1.4905 -39.6L4S Re Co.LS)t(0.5): О 03衫S. 0.25 Lo.72 Lood 0.03

1.00 3.S 11.4490 2 (o.し5x 0.05 + 0.0345) +0.01 16.0245 load (254.66 + 5L7,L3osd) = T- 134-. 9953 Xs = 0.455o·6365 ch = 62.935